The Bachelor is attempting to cross a river that

flows from north to south at 20. m/s. His top

speed is 33 m/s. At what heading should he cross

the river if he starts on the East bank and needs to

end up directly across the river on the other side? i’m


1
Expert's answer
2021-09-24T09:24:56-0400

The resultant must be directed straight from East to North. in other words, the difference between the vertical component of the Bachelor (directed North) and the river velocity (directed South) must be zero. Assume the angle (from West to North) of the Bachelor is θ:


"v_r-v_B\sin\theta=0,\\\space\\
\theta=\arcsin\frac{v_r}{v_B}=\arcsin\frac{20}{33}=37.3°."